# Marginal treatment effects in the absence of instrumental variables

### v2

@@ Abstract
We propose a method for defining, identifying, and estimating the marginal treatment effect (MTE) without imposing the instrumental variable (IV) assumptions of independence, exclusion, and separability (or monotonicity). Under a new definition of the MTE based on reduced-form treatment error that is statistically independent of the covariates, we find that the relationship between the MTE and standard treatment parameters holds in the absence of IVs. We provide a set of sufficient conditions ensuring the identification of the defined MTE in an environment of essential heterogeneity. The key conditions include a linear restriction on potential outcome regression functions, a nonlinear restriction on the propensity score, and a conditional mean independence restriction which will lead to additive separability. We prove this identification using the notion of semiparametric identification based on functional form. And we provide an empirical application for the Head Start program to illustrate the usefulness of the proposed method in analyzing heterogenous causal effects when IVs are elusive.
@@ 1 Introduction
The marginal treatment effect (MTE), which was developed in a series of seminal papers by Heckman and Vytlacil (1999, 2001, 2005), has become one of the most popular tools in social sciences for describing, interpreting, and analyzing the heterogeneity of the effect of a nonrandom treatment. The MTE is defined as the expected treatment effect conditional on the observed covariates and normalized error term consisting of unobservable determinants of the treatment status. Although the definition of MTE doesn’t necessitate an instrumental variable (IV), nearly all the existing identification strategies for the MTE rely heavily on valid instrumental variation which can induce otherwise similar individuals into different treatment choices. This is mainly because the MTE has been regarded as an extension of and supplement to the standard IV method in the causal inference literature. However, in practical applications, the validity of IVs is usually difficult to justify. In the most common case of just-identification in which only one IV is available, the exogeneity or randomness of the IV is fundamentally untestable. The exclusion restriction, which is another condition underlying the validity of IVs, has also been challenged more often than not (e.g., van den Berg, 2007; Jones, 2015).
@@
Motivated by the potential invalidity of the available instruments, and inspired by the observation that instruments are unnecessary in the definition of MTE, we model, identify, and estimate the MTE without imposing the standard IV assumptions of conditional independence, exclusion, and separability. Namely, we allow all the observed covariates to be statistically correlated to the error term and have a direct impact on the outcome in addition to an indirect impact through the treatment variable. The cornerstone of our model is a normalized treatment equation, which determines treatment participation by the propensity score crossing a reduced-form error term that is statistically independent of (although functionally dependent on) all the covariates. This is comparable to the conventional IV model, in which normalization is performed with respect to only the non-IV covariates. The independence of the reduced-form treatment error from all the covariates can partly justify a conditional mean independence assumption on the potential outcome residuals, which can ensure the additive separability of the MTE into observables and unobservables. This separability then facilitates the semiparametric identification of the MTE, given a linear restriction on the potential outcome regression functions and a nonlinear restriction on the propensity score function. We prove the identification by representing the MTE as a known function of conditional moments of observed variables. Intuitively, the identification power comes from the excluded nonlinear variation in the propensity score, which plays the role of the IV in exogenously perturbing the treatment status. Our identification strategy closely resembles the semiparametric counterpart of identification based on functional form (Dong, 2010; Escanciano et al., 2016; Lewbel, 2019; Pan et al., 2022). We build our result on the notion of identification based on functional form mainly because (i) it can lead to point identification and estimation, (ii) the required assumptions are regular, and (iii) the resulting estimation procedures are the most compatible with the standard procedures in conventional IV-MTE models. After identification, the MTE can be semiparametrically estimated by implementing the kernel-weighted pairwise difference regression for the sample selection model (Ahn and Powell, 1993) for each treatment status.
@@
We contribute to the literature on program evaluation under essential heterogeneity by proposing a method for defining, identifying, and estimating the MTE in the absence of IVs. The main value of this IV-free framework of the MTE is threefold. First, it provides an approach for consistently estimating heterogeneous causal effects when a valid IV is hard to find. Second, it suggests an easy-to-implement means for testing the validity of a candidate IV, because it nests the models that assume exclusion restrictions. For instance, in studies on returns to education, the validity of most instruments for educational attainment is suspect, such as parental education, distance to the school, and local labor market conditions (Kédagni and Mourifié, 2020; Mourifié et al., 2020). Our framework enables a reliable evaluation of returns to education without imposing IV assumptions on the candidate instruments and implies a simple test for exclusion restrictions by $t$-testing the instruments’ coefficients in potential outcome equations. Third, even though the validity of the IV is verified, identification based on functional form can be invoked as a way to increase the efficiency of estimation or to check the robustness of the results to alternative identifying assumptions.
@@
Early explorations of the identification of endogenous regression models when IVs are unavailable focused on linear systems of simultaneous equations, in which the lack of IVs is associated with hardly justifiable exclusion restrictions and insufficient moment conditions. The typical strategy for addressing this underidentification problem is to impose second- or higher-order moment restrictions to construct instruments by using the available exogenous covariates (e.g., Fisher, 1966; Lewbel, 1997; Klein and Vella, 2010). In particular, imposing restrictions on the correlation between the covariates and the variance matrix of the vector of model errors leads to identification based on heteroscedasticity (Lewbel, 2012). D’Haultfœuille et al. (2021) considered a more general nonseparable, nonparametric triangular model without the exclusion restriction, and established identification for the control function approach based on a local irrelevance condition. However, these results are restricted to the case of continuous endogenous variable or treatment. Lewbel (2018) showed that, in linear regression models, the moment conditions of identification based on heteroscedasticity can be satisfied when the endogenous treatment is binary, at the expense of strong restrictions on the model errors. Alternatively, Conley et al. (2012) and Van Kippersluis and Rietveld (2018) presented a sensitivity analysis approach for performing inference for linear IV models while relaxing the exclusion restriction to a certain extent. However, the linear models implicitly impose an undesirable homogeneity assumption on treatment effects.
@@
The literature on sample selection models without the exclusion restriction is also closely related. A general solution to the problem of lack of excluded variables is partial identification and set estimation (Honoré and Hu, 2020, 2022; Westphal et al., 2022). The limitation of this approach is that the identified or estimated set may be too wide to be informative. Another solution is the approach of identification at infinity that uses the data only for large values of a special regressor (Chamberlain, 1986; Lewbel, 2007) or of the outcome (D’Haultfœuille and Maurel, 2013). Although identification at infinity can lead to point identification, it is typically featured as irregular identification (Khan and Tamer, 2010), and the derived estimate will converge at a rate slower than $n^{-1/2}$, where $n$ is the sample size. Heckman (1979) exploited nonlinearity in the selectivity correction function to achieve point identification and root-$n$ consistent estimation for the linear coefficients of a parametric sample selection model, which is the original version of identification based on functional form. However, Heckman’s model imposes a restrictive bivariate normality assumption on the error terms and thus poses the risk of model misspecification. Escanciano et al. (2016) extended Heckman’s approach to a general semiparametric model and established identification of the linear coefficients by exploiting nonlinearity elsewhere in the model. At the expense of the generality of their model, Escanciano et al. (2016)’s identification result was only up to scale and required a scale normalization assumption. This normalization would be innocuous if we know the sign of the normalized coefficient and are interested in only the sign, but not the magnitude, of the other coefficients. However, the magnitude of the linear coefficients is indispensable to the evaluation of a program or a policy. In addition, the identifying assumption about nonlinearity developed by Escanciano et al. (2016) implicitly rules out the case of the existence of two continuous covariates. We adapt the result of Escanciano et al. (2016) to the MTE model by amending the two defects. First, we take advantage of the specific model structure to relax the scale normalization and identify the magnitude of the linear coefficients. Moreover, the identifying assumption about nonlinearity will be simplified owing to the specific model structure. We give an intuitive interpretation of the new nonlinearity assumption. Second, we combine the nonlinearity assumption with a local irrelevance condition to allow for an arbitrary number of continuous covariates.
@@
The rest of this paper is organized as follows. In Section 2, we introduce our model and the definition of MTE without IVs. In Section 3, we propose a possible set of sufficient conditions for the identification of the MTE, in place of the standard IV assumptions. The key conditions include semiparametric functional form restrictions and the conditional mean independence assumption, of which the latter implies the additive separability of the MTE into observed and unobserved components. Section 4 suggests consistent estimation procedures, and Section 5 provides an empirical application to Head Start. Section 6 concludes. Online appendices comprise (A) a proof of the primary identification result, (B) a discussion on variants of the nonlinearity assumption, and (C) an identification result for limited valued outcomes which entails a slight modification of the assumptions.
@@ 2 Model
In the following, we denote random variables or random vectors by capital letters, such as $U$, and their possible realizations by the corresponding lowercase letters, such as $u$. We denote $F_{U}\left(\cdot\right)$ as the cumulative distribution function (CDF) of $U$ and $F_{U\left|X\right.}\left(\cdot\left|x\right.\right)$ as the conditional CDF of $U$ given $X=x$. Our analysis builds on the potential outcomes framework developed by Roy (1951) in econometrics and Rubin (1973a, b) in statistics. Specifically, we consider a binary treatment, denoted by $D$, and let $Y_{1}$ and $Y_{0}$ denote the potential outcomes if the individuals are treated ($D=1$) or not treated ($D=0$), respectively. The observed outcome is
@@
Compared with the conventional MTE model, we relax the independence and separability assumptions in the treatment equation (1) to account for the absence of IVs. First, we don’t assume that $X$ is stochastically independent of $U$; that is, no exogenous covariate is needed. Second, we allow the treatment decision rule to be intrinsically nonseparable in observed and unobserved characteristics, which is equivalent to relaxing the monotonicity assumption in the IV model (Vytlacil, 2002, 2006). Specifically, let $U_{x}$ be the counterfactual error term denoting what $U$ would have been if $X$ had been externally set to $x$, then the nonseparability of (1) means that at least two vectors $x$ and $\tilde{x}$ exist in the support of $X$ such that $U_{x}\neq U_{\tilde{x}}$ with positive probability. In particular, $U$ is allowed to depend functionally on $X$, as in the following example.
@@
where the observed $X$ can be statistically correlated to the unobserved $\varepsilon$, and no restriction is imposed on the cross-partials of the index function $m$. Without independence and additive separability, model (2) is completely vacuous, imposing no restrictions on the observed or counterfactual outcomes (Heckman and Vytlacil, 2001). This general latent index rule fits into the treatment equation (1) by taking $\mu\left(X\right)=E\left[m\left(X,\varepsilon\right)\left|X\right.\right]$ and $U=\mu\left(X\right)-m\left(X,\varepsilon\right)$.
@@
$\pi\left(x\right)\equiv E\left[D\left|X=x\right.\right]=F_{U\left|X\right.}\left(\mu\left(x\right)\left|x\right.\right),$
and define the propensity score variable as $P\equiv\pi\left(X\right)$. Under the regularity condition that $F_{U\left|X\right.}\left(\cdot\left|x\right.\right)$ is absolutely continuous with respect to the Lebesgue measure for all $x$, the structural treatment equation (1) can be innocuously normalized into a reduced form:
@@
$V=F_{U\left|X\right.}\left(U\left|X\right.\right)$
is a normalized error term, which by definition follows standard uniform distribution conditional on $X$ and thus is stochastically independent of $X$ and $P$. This independence, which seems counterintuitive due to the functional dependence of $V$ on $X$, will be lost if we consider $V_{x}=F_{U\left|X\right.}\left(U_{x}\left|x\right.\right)$, the counterfactual variable of $V$ when $X$ is set to $x$. In general, the conditional distribution of $V_{x}$ given $X=\tilde{x}$ for $\tilde{x}\neq x$ depends on $\tilde{x}$, and the unconditional distribution of $V_{x}$ is not uniform.
@@
Through a property of the bivariate normal distribution, we obtain $U\left|\left(X=x\right)\right.\sim N\left(\mu_{\left.U\right|X}\left(x\right),\sigma_{U\left|X\right.}^{2}\right)$ and $F_{U\left|X\right.}\left(u\left|x\right.\right)=\Phi\left(\left.\left[u-\mu_{U\left|X\right.}\left(x\right)\right]\right/\sigma_{U\left|X\right.}\right)$, where $\mu_{U\left|X\right.}\left(x\right)=\left(\sigma_{UX}\left/\sigma_{X}^{2}\right.\right)x$, $\sigma_{U\left|X\right.}^{2}=1-\left(\sigma_{UX}^{2}\left/\sigma_{X}^{2}\right.\right)$, and $\Phi\left(\cdot\right)$ denotes the standard normal CDF. Hence,
@@
[…]
@@
$\Delta^{\text{MTE}}\left(x,v\right)$ captures all the treatment effect heterogeneity that is consequential for selection bias by conditioning on the orthogonal coordinates of the observable and unobservable dimensions. Given $X$ and $V$, the treatment status $D$ is fixed and thus independent of the treatment effect $Y_{1}-Y_{0}$. Similar to the MTE in the IV model, $\Delta^{\text{MTE}}\left(x,v\right)$ can be used as a building block for constructing the commonly used causal parameters, such as the average treatment effect (ATE), the treatment effect on the treated (TT), the treatment effect on the untreated (TUT), and the local average treatment effect (LATE), which can be expressed as weighted averages of $\Delta^{\text{MTE}}\left(x,v\right)$, as follows:
[…]
Somewhat surprisingly, compared with the weights on the conventional MTE (e.g., Heckman and Vytlacil, 2005, Table IB), which are generally difficult to estimate in practice, the weights on $\Delta^{\text{MTE}}\left(x,v\right)$ are simpler, more intuitive, and easier to compute.

### v3

@@ Abstract
We propose a method for defining, identifying, and estimating the marginal treatment effect (MTE) without imposing the instrumental variable (IV) assumptions of independence, exclusion, and separability (or monotonicity). Under a new definition of the MTE based on reduced-form treatment error that is statistically independent of the covariates, we find that the relationship between the MTE and standard treatment parameters holds in the absence of IVs. We provide a set of sufficient conditions ensuring the identification of the defined MTE in an environment of essential heterogeneity. The key conditions include a linear restriction on potential outcome regression functions, a nonlinear restriction on the propensity score, and a conditional mean independence restriction which will lead to additive separability. We prove this identification using the notion of semiparametric identification based on functional forms. And we provide an empirical application for the Head Start program to illustrate the usefulness of the proposed method in analyzing heterogenous causal effects when IVs are elusive.
@@ 1 Introduction
The marginal treatment effect (MTE), which was developed in a series of seminal papers by Heckman and Vytlacil (1999, 2001, 2005), has become one of the most popular tools in social sciences for describing, interpreting, and analyzing the heterogeneity of the effect of a nonrandom treatment. The MTE is defined as the expected treatment effect conditional on observed covariates and a normalized error term that consists of unobservable determinants of the treatment status. Although the definition of the MTE doesn’t necessitate an instrumental variable (IV), nearly all the existing identification strategies for the MTE rely heavily on valid instrumental variation which can induce otherwise similar individuals into different treatment choices. This is mainly because the MTE has been regarded as an extension of and supplement to the standard IV method in the causal inference literature. However, in practical applications, the validity of IVs is usually difficult to justify. In the most common case of just-identification in which only one IV is available, the exogeneity or randomness of the IV is fundamentally untestable. The exclusion restriction, which is another condition underlying the validity of IVs, has also been challenged more often than not (e.g., van den Berg, 2007; Jones, 2015).
@@
Motivated by the potential invalidity of the available instruments, and inspired by the observation that instruments are unnecessary in the definition of the MTE, we attempt to model, identify, and estimate the MTE without imposing the standard IV assumptions of conditional independence, exclusion, and separability in this paper. Specifically, we allow all the observed covariates to be statistically correlated to the error term and have a direct impact on the outcome in addition to an indirect impact through the treatment variable. The cornerstone of our model is a normalized treatment equation, which determines treatment participation by the propensity score crossing a reduced-form error term that is statistically independent of (although functionally dependent on) all the covariates. This is comparable to the conventional IV-based model, in which normalization is performed with respect to only non-IV covariates. The independence of the reduced-form treatment error from all the covariates can partly justify a conditional mean independence assumption on the potential outcome residuals, which can ensure the additive separability of the MTE into observables and unobservables. This separability then facilitates the semiparametric identification of MTE, given a linear restriction on the potential outcome regression functions and a nonlinear restriction on the propensity score function. We prove the identification by representing the MTE as a known function of conditional moments of observed variables. Intuitively, the identifying power comes from the excluded nonlinear variation in the propensity score, which plays the role of the IV in exogenously perturbing the treatment status.
Compared with imposing the IV assumptions, the main cost of the proposed IV-free method is the linear restriction imposed on the outcome equations and the nonlinearity imposed on the treatment equation, which makes our identification strategy for the MTE closely resemble the semiparametric counterpart of identification based on functional forms (Dong, 2010; Escanciano et al., 2016; Lewbel, 2019; Pan et al., 2022). We build our result on the notion of identification based on functional forms mainly because (i) it can lead to point identification and estimation, (ii) the required assumptions are regular and partly testable, and (iii) the resulting estimation procedures are the most compatible with the standard procedures in IV-based MTE models. Although the primary contribution of this work is on the identification of MTE in the absence of IVs, we also propose a semiparametric MTE estimator tailored to the identification result by implementing the kernel-weighted pairwise difference regression (Ahn and Powell, 1993) for each treatment status, and establish its consistency and asymptotic normality that take into account the nonparametrically estimated propensity score in the first step.
@@
The main value of our IV-free framework of the MTE is threefold. First, it provides an approach for consistently estimating heterogeneous causal effects when a valid IV is hard to find. Second, it suggests an easy-to-implement means for testing the validity of a candidate IV, because it nests the models that assume exclusion restrictions. For instance, in studies on returns to education, the validity of most instruments for educational attainment is suspect, such as parental education, distance to the school, and local labor market conditions (Kédagni and Mourifié, 2020; Mourifié et al., 2020). Our framework enables a reliable evaluation of returns to education without imposing IV assumptions on the candidate instruments and implies a simple test for exclusion restrictions by $t$-testing the instruments’ coefficients in potential outcome equations. Third, even though the validity of the IV is verified, identification based on functional forms can be invoked as a way to increase the efficiency of estimation or to check the robustness of the results to alternative identifying assumptions.
@@
Early explorations of the identification of endogenous regression models when IVs are unavailable focused on linear systems of simultaneous equations, in which the lack of IVs is associated with hardly justifiable exclusion restrictions and insufficient moment conditions. The typical strategy for addressing this underidentification problem is to impose second- or higher-order moment restrictions to construct instruments by using the available exogenous covariates (e.g., Fisher, 1966; Lewbel, 1997; Klein and Vella, 2010). In particular, imposing restrictions on the correlation between the covariates and the variance matrix of the vector of model errors leads to identification based on heteroscedasticity (Lewbel, 2012). D’Haultfœuille et al. (2024) considered a more general nonseparable, nonparametric triangular model without the exclusion restriction, and established identification for the control function approach based on a local irrelevance condition. However, these results are restricted to the case of continuous endogenous variable or treatment. Lewbel (2018) showed that, in linear regression models, the moment conditions of identification based on heteroscedasticity can be satisfied when the endogenous treatment is binary, at the expense of strong restrictions on the model errors. Alternatively, Conley et al. (2012) and Van Kippersluis and Rietveld (2018) presented a sensitivity analysis approach for performing inference for linear IV models while relaxing the exclusion restriction to a certain extent. Kang et al. (2016) proposed an identification and estimation method for linear treatment effect models when the exclusion restriction of some IVs may not hold, and Windmeijer et al. (2019) further proposed a robust method to detect invalid IVs that enter the outcome equation as explanatory variables. More recently, Masten and Poirier (2021) suggested a set identification method for linear IV models under violations of the exclusion restriction. A common limitation of this body of literature is that linear models implicitly impose an undesirable homogeneity assumption on treatment effects. Carl et al. (2025) considered invalid IVs in a treatment effect model with nonlinearity in both the treatment and outcome equations, and established the identification based on different functional forms in the two equations, which is similar to our identification strategy in terms of ideas. Moreover, they adopted machine learning methods in estimating the treatment equation to cope with the potentially high nonlinearity. However, their model implicitly assumes homogenous treatment effects as well. As a complement, our model relaxes the restrictive homogeneity assumption and allows for general heterogeneities in treatment effects, which is evidently more realistic.
@@
The literature on sample selection models without the exclusion restriction is also closely related. A general solution to the problem of lack of excluded variables is partial identification and set estimation (Honoré and Hu, 2020, 2024; Westphal et al., 2022). The limitation of this approach is that the identified or estimated set may be too wide to be informative. Another solution is the approach of identification at infinity that uses the data only for large values of a special regressor (Chamberlain, 1986; Lewbel, 2007) or of the outcome (D’Haultfœuille and Maurel, 2013). Although identification at infinity can lead to point identification, it is typically featured as irregular identification (Khan and Tamer, 2010), and the derived estimate will converge at a rate slower than $n^{-1/2}$, where $n$ is the sample size. Heckman (1979) exploited nonlinearity in the selectivity correction function to achieve point identification and root-$n$ consistent estimation for the linear coefficients of a parametric sample selection model, which is the original version of identification based on functional forms. However, Heckman’s model imposes a restrictive bivariate normality assumption on the error terms and thus poses the risk of model misspecification. Escanciano et al. (2016) extended Heckman’s approach to a general semiparametric model and established identification of the linear coefficients by exploiting nonlinearity elsewhere in the model. At the expense of the generality of their model, Escanciano et al. (2016)’s identification result is only up to scale and thus requires a scale normalization assumption. This normalization would be innocuous if we know the sign of the normalized coefficient and are interested in only the sign, but not the magnitude, of the other coefficients. However, the magnitude of the linear coefficients is indispensable to the evaluation of a program or a policy. In addition, the identifying assumption about nonlinearity developed by Escanciano et al. (2016) implicitly rules out the case of the existence of two continuous covariates. We adapt the result of Escanciano et al. (2016) to the MTE model by amending the two defects. First, we take advantage of the specific model structure to relax the scale normalization and identify the magnitude of the linear coefficients. Moreover, the identifying assumption about nonlinearity will be simplified owing to the specific model structure. We give an intuitive interpretation of the new nonlinearity assumption. Second, we combine the nonlinearity assumption with a local irrelevance condition to allow for an arbitrary number of continuous covariates.
@@
The rest of this paper is organized as follows. In Section 2, we introduce our model and definition of the MTE without IVs. In Section 3, we propose a possible set of sufficient conditions for the identification of MTE, in place of the standard IV assumptions. The key conditions include semiparametric functional form restrictions and the conditional mean independence assumption, of which the latter implies the additive separability of the MTE into observed and unobserved components. Section 4 suggests consistent estimation procedures, and Section 5 provides an empirical application to Head Start. Section 6 concludes. Online appendices comprise of (A) additional results of the empirical application, (B) a proof of the main identification result, (C) a discussion on variants of the nonlinearity assumption, (D) an identification result for limited valued outcomes which entails a slight modification of the assumptions, (E) a proof of the asymptotic normality of the semiparametric IV-free MTE estimator, (F) an implementation guidance, and (G) a simulation study.
@@ 2 Model
[…]
@@
Compared with the IV-based MTE model, we relax the independence and separability assumptions in the treatment equation (1) to account for the absence of IVs. First, we don’t assume that $X$ is stochastically independent of $U$; that is, no exogenous covariate is needed. Second, we allow the treatment decision rule to be intrinsically nonseparable in observed and unobserved characteristics, which is equivalent to relaxing the monotonicity assumption in the IV model (Vytlacil, 2002, 2006). Specifically, let $U_{x}$ be the counterfactual error term denoting what $U$ would have been if $X$ had been externally set to $x$, then the nonseparability of (1) means that at least two vectors $x$ and $\tilde{x}$ exist in the support of $X$ such that $U_{x}\neq U_{\tilde{x}}$ with positive probability. In particular, $U$ is allowed to depend functionally on $X$, as in the following example.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
We observe that $V\perp\!\!\!\!\perp X$ because $F_{V\left|X\right.}\left(v\left|x\right.\right)=v$, but $V_{x}$ is not independent of $X$ because
@@
$\Delta^{\text{MTE}}\left(x,v\right)$ captures all the treatment effect heterogeneity that is consequential for selection bias by conditioning on the orthogonal coordinates of the observable and unobservable dimensions. Given $X$ and $V$, the treatment status $D$ is fixed and thus independent of the treatment effect $Y_{1}-Y_{0}$. Note that if the assumptions of independence and separability hold for some covariates, the normalized error term $V$ defined in (4) will be exactly the same as that in the IV-based MTE model, and if further the independent covariates have no direct effect on the potential outcomes, our defined MTE will reduce to the MTE that is defined in the IV model. Similar to the MTE in the IV model, the IV-free $\Delta^{\text{MTE}}\left(x,v\right)$ can be used as a building block for constructing the commonly used causal parameters, such as the average treatment effect (ATE), the treatment effect on the treated (TT), the treatment effect on the untreated (TUT), and the local average treatment effect (LATE), which can be expressed as weighted averages of $\Delta^{\text{MTE}}\left(x,v\right)$ as follows:
$\Delta^{\text{ATE}}\left(x\right)$ $\equiv$ $E\left[\left.Y_{1}-Y_{0}\right|X=x\right]=\int_{0}^{1}\Delta^{\text{MTE}}\left(x,v\right)dv,$
$\Delta^{\text{TT}}\left(x\right)$ $\equiv$ $E\left[\left.Y_{1}-Y_{0}\right|X=x,D=1\right]=\frac{1}{\pi\left(x\right)}\int_{0}^{\pi\left(x\right)}\Delta^{\text{MTE}}\left(x,v\right)dv,$
$\Delta^{\text{TUT}}\left(x\right)$ $\equiv$ $E\left[\left.Y_{1}-Y_{0}\right|X=x,D=0\right]=\frac{1}{1-\pi\left(x\right)}\int_{\pi\left(x\right)}^{1}\Delta^{\text{MTE}}\left(x,v\right)dv,$
$\Delta^{\text{LATE}}\left(x,v_{1},v_{2}\right)$ $\equiv$ $E\left[\left.Y_{1}-Y_{0}\right|X=x,v_{1}\leq V\leq v_{2}\right]=\frac{1}{v_{2}-v_{1}}\int_{v_{1}}^{v_{2}}\Delta^{\text{MTE}}\left(x,v\right)dv.$
Somewhat surprisingly, compared with the weights on the IV-based MTE (e.g., Heckman and Vytlacil, 2005, Table IB), which are generally difficult to estimate in practice, the weights on $\Delta^{\text{MTE}}\left(x,v\right)$ are simpler, more intuitive, and easier to compute.